a 12-kg block is acted upon by two forces that are in opposite direction, one a 540 Newton and the other of 1260 N. Determine the acceleration produced. How is the direction of the acceleration related to the forces?


1
Expert's answer
2021-02-28T07:36:15-0500

Let the first force of 540 N acts upon a block in the positive "x"-direction, and the second one - 1260 N acts in the opposite direction. Let's find the net force that acts upon the block:


"F_{net}=F_1-F_2=540\ N-1260\ N=-720\ N."

The sign minus means that the net force that acts upon the block directed in the opposite direction.

Let's find the acceleration from the Newton's Second Law of Motion:


"a=\dfrac{-F_{net}}{m}=\dfrac{-720\ N}{12\ kg}=-60\ \dfrac{m}{s^2}."

The sign minus means that the block accelerates under the action of the net force in the opposite direction (-"x"-direction).